HETAV PATEL

Grade 11,

The solubility product constant of Ca(OH)2 at25 degree celcius is 4.42*10^-5.A 500Ml of saturated solution of Ca(OH)2 is mixed with equal volume of 0.4M NaOH. What is mass of Ca(OH)2 precipitate out?

The solubility product constant of Ca(OH)2 at25 degree celcius is 4.42*10^-5.A 500Ml of saturated solution of Ca(OH)2 is mixed with equal volume of 0.4M NaOH. What is mass of Ca(OH)2 precipitate out?
 
 
 

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
From the question:
ionisation eqn. of Ca(Oh)2: Ca(OH)2 = Ca2+ + 2OH–
ksp of Ca(OH)2=[Ca 2+] [OH-]2=s*(2s)^2=4s^3
=>s=cube-root(ksp/4)=sqrt(4.42*10^-5 /4)=2.23*10^-2 M
Ca 2+ ions in 500 ml  soln.=(2.23*10^-2)/2 =0.011
500 ml soln+500 ml NaOH=1000ml,molarity of OH- in new solution=0.2M
[Ca]2+ =ksp/[OH-]^2 = (4.42*10^-5)/(0.2*0.2)=0.0011M
no. of moles of Ca2+ ppt=0.011-0.0011=0.0102M. Mass of Ca(OH)2= 0.0102*74.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)